Calculate the 4 point moving averages for this list of data 5,3,6,8,7,5,3,5.


1
Expert's answer
2021-01-12T14:45:17-0500

"MA_i=\frac{x_i+x_{i+1}+x_{i+2}+x_{i+3}}4\\
Hence\\
MA_1=\frac{5+3+6+8}4=5.5\\
MA_2=6\\
MA_3=6.5\\
MA_4=5.75\\
MA_5=5"